# Triplet scalar flavored leptogenesis with spontaneous CP violation

### v1

@@ 4.4 Summary
    The hierarchical branching ratios result in different amounts of washout effects along different flavor directions. The amount of washouts greatly influences the efficiency of leptogenesis. Comparing figures (3(b)) ($B_{1}^{L}=0.58$) and (4(b)) ($B_{2}^{L}=0.0006$), one can see that in the first case, the washout effects along the flavor directions are more as compared to the second case (except $ee$-flavor direction). This feature can be attributed to the specific flavor structure of the triplet Yukawa coupling matrices of the model. Although, in the second case, the extreme hierarchy in the branching ratios can in principle suppress the flavored CP asymmetry parameters, the presence of SCPV phase $\beta$ can neutralise the effect.
@@
      From Fig. (6), for both cases - I and II, with hierarchical branching ratios the baryon asymmetry increases for a fixed value of $M_{1}(M_{2})$. As shown in tables (3) and (5), the suppression in CP asymmetry due to the hierarchical branching ratio, can be compensated by the extra source of CP violation due to the presence of the SCPV phase $\beta$ in the neutrino mass matrix. As a result, even with the modified washout effect from hierarchical branching ratios, it is possible to achieve maximal CP violation and baryon asymmetry can be achieved in the observable range. The flavor structure of the model is suitable to draw these conclusions.
@@ 5 Conclusion
Spontaneous violation of CP symmetry is very promising in explaining CP violation in low-energy processes like neutrino oscillation and high-energy process like leptogenesis. In general CP violation in these two sectors are not related [21] and are model-dependent. In this paper, we have studied, in an extended Standard model, neutrino mass generated via type-II seesaw mechanism and leptogenesis from the decay of heavy scalar triplet with a common source of CP violation. The model provides a viable explanation of baryon asymmetry through thermal leptogenesis, by sourcing enough CP asymmetry spontaneously broken due to the presence of a scalar singlet with complex VEV. Moreover, the neutrino mass matrix exhibits TBM structure which generally fails to address the occurrence of non-zero CHOOZ mixing angle, $\theta_{13}$. By incorporating small perturbations around flavon VEV in the high-energy scale, it can accommodate $\theta_{13}\neq 0$ in the model. With the latest neutrino oscillation data, bound on total neutrino mass and the effective neutrino mass, we have analyzed the parameter space of the model. Using the constrained parameter space of the model we have studied flavored leptogenesis from the decay of the triplet scalars. The flavor structure of the model suppresses flavor-independent CP asymmetry and hence we explore the effect of flavor in leptogenesis by using a set of flavor-covariant Boltzmann equations considering a density matrix formalism. In our study, the scale of leptogenesis is $\mathcal{O}(10^{10})$ GeV. We also study the role of hierarchical branching ratios and the SCPV phase $\beta$ in generating maximal CP asymmetry to generate baryon asymmetry in the observational range. Requiring mass of triplet as low as possible, we find BAU in the observable range can be produced for $M_{1}=3\times 10^{10}$ GeV for comparable branching ratios. Whereas the same can be achieved for $M_{2}=1.4\times 10^{10}$ GeV for hierarchical branching ratios. The flavor structure of the triplet Yukawa couplings plays an important role in generating the washout effects which are inevitable for thermal leptogenesis.
@@ Appendix C Estimation of flavored CP asymmetries
$\epsilon^{ee}_{1}=\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\Big{[}\frac{2}{3}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\left(1-\varepsilon_{1}-\varepsilon_{2}\right)\sin(\beta-\sigma_{1})\Big{]},$ (C.8)
$\epsilon^{e\mu}_{1}$$\approx$ $\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\Big{[}-\frac{2}{3}\varepsilon_{1}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})\Big{]}$ (C.9)
@@
$\epsilon^{\mu\mu}_{1}$$\approx$ $\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\times\varepsilon_{2}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.10)
@@
$\epsilon^{\tau\tau}_{1}$$\approx$ $\frac{M^{2}_{1}|u_{1}|^{2}}{2\pi\left(3z^{2}_{1}v^{4}+4M^{4}_{1}|u_{1}|^{4}\right)}\times\varepsilon_{1}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.11)
@@
$\epsilon^{ee}_{2}$ $=$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times-\frac{2}{3}z_{2}$ (C.12)
@@
$\epsilon^{e\mu}_{2}$$\approx$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times\frac{1}{3}z_{2}$ (C.13)
@@
$\epsilon^{\mu\mu}_{2}$$\approx$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times-\frac{2}{3}z_{2}$ (C.14)
@@
$\epsilon^{\tau\tau}_{2}$$\approx$ $\frac{M^{2}_{2}|u_{2}|^{2}}{4\pi\left(z^{2}_{2}v^{4}(1-2\varepsilon_{1}-2\varepsilon_{2})+2M^{4}_{2}|u_{2}|^{4}\right)}\times-\frac{2}{3}z_{2}$ (C.15)

### v2

@@ 4.4 Summary
    The hierarchical branching ratios result in different amounts of washout effects along different flavor directions. The amount of washouts greatly influences the efficiency of leptogenesis. Comparing figures (4(b)) ($B_{1}^{L}=0.58$) and (5(b)) ($B_{2}^{L}=0.0006$), one can see that in the first case, the washout effects along the flavor directions are more as compared to the second case (except $ee$-flavor direction). This feature can be attributed to the specific flavor structure of the triplet Yukawa coupling matrices of the model. Although, in the second case, the extreme hierarchy in the branching ratios can in principle suppress the flavored CP asymmetry parameters, the presence of SCPV phase $\beta$ can neutralise the effect.
@@
      From Fig. (6), for both cases - I and II, with hierarchical branching ratios the baryon asymmetry increases for a fixed value of $M_{1}(M_{2})$. As shown in tables (3) and (5), the suppression in CP asymmetry due to the hierarchical branching ratio, can be compensated by the extra source of CP violation due to the presence of the SCPV phase $\beta$ in the neutrino mass matrix. As a result, even with the modified washout effect from hierarchical branching ratios, it is possible to achieve maximal CP violation and baryon asymmetry in the observable range. The flavor structure of the model is suitable to draw these conclusions.
@@ 5 Conclusion
Spontaneous violation of CP symmetry is very promising in explaining CP violation in low-energy processes like neutrino oscillation and high-energy process like leptogenesis. In general CP violation in these two sectors are not related [21] and are model-dependent. In this paper, we have studied, in an extended Standard model, neutrino mass generated via type-II seesaw mechanism and leptogenesis from the decay of heavy scalar triplet with a common source of CP violation. The model provides a viable explanation of baryon asymmetry through thermal leptogenesis, by sourcing enough CP symmetry spontaneously broken due to the presence of a scalar singlet with complex VEV. Moreover, the neutrino mass matrix exhibits TBM structure which generally fails to address the occurrence of non-zero CHOOZ mixing angle, $\theta_{13}$. By incorporating small perturbations around flavon VEV in the high-energy scale, it can accommodate $\theta_{13}\neq 0$ in the model. With the latest neutrino oscillation data and bound on the effective neutrino mass, we have analyzed the parameter space of the model. Using the constrained parameter space of the model we have studied flavored leptogenesis from the decay of the triplet scalars. The flavor structure of the model suppresses flavor-independent CP asymmetry and hence we explore the effect of flavor in leptogenesis by using a set of flavor-covariant Boltzmann equations considering a density matrix formalism. In our study, the scale of leptogenesis is $\mathcal{O}(10^{10})$ GeV. We also study the role of hierarchical branching ratios and the SCPV phase $\beta$ in generating maximal CP asymmetry to generate baryon asymmetry in the observational range. Requiring mass of triplet as low as possible, we find BAU in the observable range can be produced for $M_{1}=3\times 10^{10}$ GeV for comparable branching ratios. Whereas the same can be achieved for $M_{2}=1.4\times 10^{10}$ GeV for hierarchical branching ratios. The flavor structure of the triplet Yukawa couplings plays an important role in generating the washout effects which are inevitable for thermal leptogenesis.
@@ Appendix C Estimation of flavored CP asymmetries
$\epsilon^{ee}_{1}=\frac{M_{1}(B^{L}_{1}B^{\phi}_{1})^{\frac{1}{2}}c_{ee}}{v^{2}[{\rm Tr}(m^{(1)\dagger}_{\nu}m^{(1)}_{\nu})]^{1/2}}\Big{[}\frac{2}{3}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\left(1-\varepsilon_{1}-\varepsilon_{2}\right)\sin(\beta-\sigma_{1})\Big{]},$ (C.8)
$\epsilon^{e\mu}_{1}$$\approx$ $\frac{M_{1}(B^{L}_{1}B^{\phi}_{1})^{\frac{1}{2}}c_{e\mu}}{v^{2}[{\rm Tr}(m^{(1)\dagger}_{\nu}m^{(1)}_{\nu})]^{1/2}}\Big{[}-\frac{2}{3}\varepsilon_{1}z_{1}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})\Big{]}$ (C.9)
@@
$\epsilon^{\mu\mu}_{1}$$\approx$ $\frac{M_{1}(B^{L}_{1}B^{\phi}_{1})^{\frac{1}{2}}c_{\mu\mu}}{v^{2}[{\rm Tr}(m^{(1)\dagger}_{\nu}m^{(1)}_{\nu})]^{1/2}}\times\varepsilon_{2}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.10)
@@
$\epsilon^{\tau\tau}_{1}$$\approx$ $\frac{M_{1}(B^{L}_{1}B^{\phi}_{1})^{\frac{1}{2}}c_{\tau\tau}}{v^{2}[{\rm Tr}(m^{(1)\dagger}_{\nu}m^{(1)}_{\nu})]^{1/2}}\times\varepsilon_{1}z_{1}\Big{[}\frac{1}{3}\left(z^{2}_{1}+z^{2}_{2}+2z_{1}z_{2}\cos\beta\right)^{\frac{1}{2}}\sin(\beta-\sigma_{1})$ (C.11)
@@
$\epsilon^{ee}_{2}$ $=$ $\frac{M_{2}(B^{L}_{2}B^{\phi}_{2})^{\frac{1}{2}}c_{ee}}{v^{2}[{\rm Tr}(m^{(2)\dagger}_{\nu}m^{(2)}_{\nu})]^{1/2}}\times-\frac{2}{3}z_{2}\times$ (C.12)
@@
$\epsilon^{e\mu}_{2}$$\approx$ $\frac{M_{2}(B^{L}_{2}B^{\phi}_{2})^{\frac{1}{2}}c_{e\mu}}{v^{2}[{\rm Tr}(m^{(2)\dagger}_{\nu}m^{(2)}_{\nu})]^{1/2}}\times\frac{1}{3}z_{2}\times$ (C.13)
@@
$\epsilon^{\mu\mu}_{2}$$\approx$ $\frac{M_{2}(B^{L}_{2}B^{\phi}_{2})^{\frac{1}{2}}c_{\mu\mu}}{v^{2}[{\rm Tr}(m^{(2)\dagger}_{\nu}m^{(2)}_{\nu})]^{1/2}}\times-\frac{2}{3}z_{2}\times$ (C.14)
@@
$\epsilon^{\tau\tau}_{2}$$\approx$ $\frac{M_{2}(B^{L}_{2}B^{\phi}_{2})^{\frac{1}{2}}c_{\tau\tau}}{v^{2}[{\rm Tr}(m^{(2)\dagger}_{\nu}m^{(2)}_{\nu})]^{1/2}}\times-\frac{2}{3}z_{2}\times$ (C.15)